# WiNGPT-3.0 Technical Report

### v1

@@ Abstract
Current Large Language Models (LLMs) exhibit significant limitations, notably in structured, interpretable, and verifiable medical reasoning, alongside practical deployment challenges related to computational resources and data privacy. This report focused on the development of WiNGPT-3.0, the 32-billion parameter LLMs, engineered with the objective of enhancing its capacity for medical reasoning and exploring its potential for effective integration within healthcare IT infrastructures. The broader aim is to advance towards clinically applicable models. The approach involved a multi-stage training pipeline tailored for general, medical, and clinical reasoning. This pipeline incorporated supervised fine-tuning (SFT) and reinforcement learning (RL), leveraging curated Long Chain-of-Thought (CoT) datasets, auxiliary reward models, and an evidence-based diagnostic chain simulation. WiNGPT-3.0 demonstrated strong performance: specific model variants achieved scores of 66.6 on MedCalc and 87.1 on MedQA-USMLE. Furthermore, targeted training improved performance on a clinical reasoning task from a baseline score of 58.1 to 62.5. These findings suggest that reinforcement learning, even when applied with a limited dataset of only a few thousand examples, can enhance medical reasoning accuracy. Crucially, this demonstration of RL’s efficacy with limited data and computation paves the way for more trustworthy and practically deployable LLMs within clinical workflows and health information infrastructures.
@@ 6 Discussion
RL markedly improved alignment with clinical-reasoning principles. WiNGPT-3.0-S3 outperformed the base Qwen-32B on MedCalc (66.6 vs. 51.4) and MedReMCQ (74.7 vs. 73.5), indicating stronger medical reasoning ability. Merging with a strong base model (QwQ-32B) prior to RL fine-tuning (WiNGPT-3.0-S3-Merged) yielded substantial benefits; its superior performance across diverse benchmarks—medical (MedQA-USMLE: 87.1; MedMCQA: 71.6), general (CMMLU: 86.1), and mathematical reasoning (MATH-500: 95.6)—indicates that this integration effectively preserves general capabilities while facilitating domain-specific adaptation. Nonetheless, current benchmarks still under-represent real-world complexity: they emphasise pattern matching, overlook temporal factors, and offer limited probes for step-wise reasoning.
@@ 7 Conclusion
Future research will concentrate on developing advanced training methodologies for complex medical scenarios, improving model reliability, safety, and interpretability, exploring multi-modal data integration, and rigorously evaluating its potential to deliver positive societal impact, particularly within the healthcare sector.

### v2

@@ Abstract
Current Large Language Models (LLMs) exhibit significant limitations, notably in structured, interpretable, and verifiable medical reasoning, alongside practical deployment challenges related to computational resources and data privacy. This report focused on the development of WiNGPT-3.0, a 32-billion parameter LLM, engineered with the objective of enhancing its capacity for medical reasoning and exploring its potential for effective integration within healthcare IT infrastructures. The broader aim is to advance towards clinically applicable models. The approach involved a multi-stage training pipeline tailored for general, medical, and clinical reasoning. This pipeline incorporated supervised fine-tuning (SFT) and reinforcement learning (RL), leveraging curated Long Chain-of-Thought (CoT) datasets, auxiliary reward models, and evidence-based diagnostic chain simulation. WiNGPT-3.0 demonstrated strong performance: specific model variants achieved scores of 66.6 on MedCalc and 87.1 on MedQA-USMLE. Furthermore, targeted training improved performance on a clinical reasoning task from a baseline score of 58.1 to 62.5. These findings suggest that reinforcement learning, even when applied with a limited dataset of only a few thousand examples, can improve the accuracy of medical reasoning. Crucially, this demonstration of RL’s efficacy with limited data and computation paves the way for more trustworthy and practically deployable LLMs within clinical workflows and health information infrastructures. WiNGPT-3.0 visit information, demos and updates at https://github.com/winninghealth/WiNGPT-3.0.
@@ 6 Discussion
RL markedly improved alignment with clinical-reasoning principles. WiNGPT-3.0-S3 outperformed the base QwQ-32B on MedCalc (66.6 vs. 51.4) and MedReMCQ (74.7 vs. 73.5), indicating stronger medical reasoning ability. Merging with a strong model (QwQ-32B) prior to RL fine-tuning (WiNGPT-3.0-S3-Merged) yielded substantial benefits; its superior performance across diverse benchmarks—medical (MedQA-USMLE: 87.1; MedMCQA: 71.6), general (CMMLU: 86.1), and mathematical reasoning (MATH-500: 95.6)—indicates that this integration effectively preserves general capabilities while facilitating domain-specific adaptation. Nonetheless, current benchmarks still under-represent real-world complexity: they emphasise pattern matching, overlook temporal factors, and offer limited probes for step-wise reasoning.
@@ 7 Conclusion
Future research will concentrate on developing advanced training methodologies for complex medical scenarios, improving model reliability, safety, and interpretability, exploring multi-modal data integration, and rigorously evaluating its potential to deliver positive societal impact, particularly within the healthcare sector. We aim to further reduce the cost of running LLMs, enabling their wider application in real-world scenarios.